OpenStudy (anonymous):

an=(1/2)^n-1 start with n=1 and first 5 terms

OpenStudy (anonymous):

a1=(1/2)^1-1 =(1/2)^0 a1 =1 a2=(1/2)^2-1 =(1/2)^1 a2=1/2 d =( 1/2)-1 d = (-1/2) a5= 1+ (5-1)*(-1/2) = 1+4^(-1/2) = 1+(1/(2^2(1/2)) = 1+(1/2) a5=3/2